OpenStudy (anonymous):

Which of the following equations have the same solution? Solve all three equations separately and explain each step. I. 2x + 4 = 3x – 11 II. 3x + 3 = 51 III. 2x+1/3=x+6/2

OpenStudy (anonymous):

@KingGeorge help me on this one?

OpenStudy (anonymous):

@surjithayer

OpenStudy (anonymous):

Solve each for x.

OpenStudy (anonymous):

2x+4=3x-11 -2x -2x 4=x-11 +11 +11 x=15 ?

OpenStudy (anonymous):

3x + 3 = 51 -3 -3 3x= 48 x=16 ?

OpenStudy (anonymous):

i solve III \[\frac{2x+ 1 }{ 3 }=\frac{ x+6 }{ 2 }\] cross multiply 4x+2=3x+18 4x-3x=18-2 x=16

OpenStudy (anonymous):

o.o Yuh

OpenStudy (anonymous):

ii correct

OpenStudy (anonymous):

i correct.

OpenStudy (anonymous):

thanks @surjithayer im taking my test and im having a hard time

OpenStudy (anonymous):

you see 2nd and 3rd equations have same solution. you are welcome.

OpenStudy (anonymous):

do yu mind helping me with a few questions? i dont want a direct answer i just want yu to guide me while i try to figure out the answer,

OpenStudy (anonymous):

@surjithayer

OpenStudy (anonymous):

sure i will try.

OpenStudy (anonymous):

Neveah has created the function f(x) = 2x +4/3 to represent the growth of her hair, where x represents the number of months since her last hair cut. Neveah discovers that using the inverse function to solve for x = 6, she can predict when her hair will have grown 6 inches. Explain to Neveah how to accomplish this, using complete sentences.